Aniline, C6H5NH2, when reacted with picric acid gives a derivative with an absorptivity of 

 134 cm−1

g

−1 L at 359 nm. What would be the absorbance of a 1.00 × 10−4 M solution of 

 reacted aniline in a 1.00-cm cell


1
Expert's answer
2021-08-16T02:46:07-0400

Solution:

The molar mass of aniline is 93.13 g mol−1.

Therefore,

(134 cm−1 g−1 L) × (93.13 g mol−1) = 12479.42 cm−1 L mol−1 = 12479.42 cm−1 M-1

ε = 12479.42 cm−1 M-1


The Beer–Lambert law can be used.

A = ε × l × C

where:

A = absorbance (no units)

ε = molar absorptivity (cm−1 M-1)

l = the pathlength (cm)

C = concentration


Therefore,

A = (12479.42 cm−1 M-1) × (1.00 cm) × (1.00 × 10−4 M) = 1.2479 = 1.25

A = 1.25


Answer: The absorbance (A) of a solution is 1.25